Suppose that Lil John Industries’ equity is currently selling for $29 per share and that 2.2 million shares are outstanding. The firm also has 52,000 bonds outstanding, which are selling at 104 percent of par. Assume Lil John was considering an active change to its capital structure so that the firm would have a (D/E) of 1.3.

Which type of security (stocks or bonds) would it need to sell to accomplish this?

  • sell bonds and buy back stock

  • sell stocks and buy back bonds

How much would the firm have to sell? (Enter your answer in dollars not in millions. Do not round intermediate calculations and round your final answer to 2 decimal places.)
Answer #1

Current value of equity= Number of shares* price

=2,200,000* 29

=$63,800,000.00

Current value of bonds = 52000*104%*1000=54080,000

Total value = Equity+ Bonds

= 63800,000+ 54,080,000

=117880000

D/E ratio = 1.3/1

Total value= (Debt+Equity)

(1.3+1) = 2.3

Debt = 1.3*117880000/2.3=66627826.09

Equity = 1*117880000/2.3 = 51252173.91

Hence the firm needs to sell equity and issue bonds

Equity to be sold = 63800000- 51252173.91

=$12547826.09